Consider the temperature-vapor pressure data for ammonia:

at -63.15 celcius, vapor pressure= 134.3 torr

at -38.15 celcius, vapor pressure= 597.0 torr.

determine the normal boiling point for ammonia.
Answer #1

temperature T1 = - 63.15 oC = 210 K

T2 = - 38.15 oC = 235 K

pressure P1 = 134.3 torr

P2 = 597.0 torr

ln (p2 / p1) = delta H / R [1 / T1 - 1/ T2]

ln (597 / 134.3) = delta H / 8.314 x 10^-3 [1/210 - 1/235]

delta H = 24.48 kJ/mol

pressure of ammonia = 760 torr

ln (760 / 597) = 24.48 / 8.314 x 10^-3 [1/235 - 1/ T2]

T2 = 239.6 K = - 33.5 oC

normal boiling point of ammonia = - 33.53 oC